# Sparse Autoencoder Features for Classifications and Transferability

### v1

@@ 1 Introduction
Sparse Autoencoders (SAEs) have recently emerged as a promising technique within MI for understanding LLMs. SAEs generally work by learning a compressed, sparse representation of the LLM’s internal activations. This is achieved by up-projecting the dense hidden state of the LLM to a sparser, ideally monosemantic, representation Bricken et al. (2023); Cunningham et al. (2023). Identifying semantically meaningful features within LLMs using SAEs allows for deploying these features into explainable classification pipelines. This has the potential to boost performance and detect harmful biases or spurious correlations before they manifest in downstream tasks Anthropic Interpretability Team (2024). The ability to employ SAE features for classification across diverse settings, ranging from toxicity detection to user intent, offers a scalable form of "model insight" Bowman et al. (2022), which is crucial for building trust, safety, and accountability in high-stakes domains like medicine and law Abdulaal et al. (2024).
Despite the promise of SAEs for MI, surprisingly few systematic studies have provided practical guidance on their use for classification. While promising results have been reported across various tasks Elhage et al. (2024); Kantamneni et al. (2024); Chen et al. (2024), inconsistent experimental protocols, a lack of standardized benchmarks, and limited exploration of key architectural decisions hinder comparability and the development of best practices. Although tools like Transformer Lens Nanda and Bloom (2022) and SAE Lens Joseph Bloom and Chanin (2024) have improved standardization in sampling activations, critical questions about optimal configurations for diverse tasks, particularly in multilingual and multimodal settings, remain unanswered. This makes it challenging to establish the robustness and generalizability of SAE-based classification approaches.
@@ 2.1 Interpretable Feature Extraction
SAEs address these limitations by providing more interpretable feature sets Bricken et al. (2023); Cunningham et al. (2023). Recent scaling efforts have demonstrated SAE viability across LLMs from Claude 3 Sonnet Templeton et al. (2024) to GPT-4 Gao et al. (2024) with extensions to multimodal architectures like CLIP Bhalla et al. (2024). Although these studies have revealed interpretable feature dimensions and computational circuits Marks et al. (2024); Zhao et al. (2024), they focus mainly on descriptive feature discovery rather than systematic evaluation of their downstream applications. Our work bridges this gap by providing standardized evaluation frameworks for SAE-based classification and cross-modal transfer, establishing quantitative metrics and methods for feature utility across diverse tasks.
@@ 2.2 SAE-Based Classification and its Limitations
Reports have demonstrated that SAE-derived features can outperform traditional hidden-state probing for classification, particularly in scenarios with noisy or limited data with closed datasets Anthropic Interpretability Team (2024) or simplified tasks Kantamneni et al. (2024). However, more recent studies, such as Wu et al. (2025), suggest that SAEs may not be superior, particularly for model steering (instead of classification). These seemingly conflicting results highlight a critical gap in the current understanding of SAE-based classification: a lack of systematic exploration of how hyperparameters, feature aggregation strategies, and other methodological choices impact performance.
@@ 3 Preliminaries
Since SAEs generate token-level feature activations, an essential step in classification is aggregating these activations into a fixed-size sequence representation. Without pooling, the model lacks a structured way to combine token-level representations. Previous NLP works have explored various pooling strategies for feature aggregation in neural representations Shen et al. (2018). However, it remains unclear which pooling method is most effective for LLMs’ SAE features. We systematically evaluate different pooling approaches (displayed in 2, considering (1) Top-$N$ feature selection per token  [^3] Token-level top-$N$ where n=0 indicates the absence of max pooling. and (2) summation-based aggregation444this approach is also adopted by parallel research Brinkmann et al. (2025). which collapses token-level activations into a single sequence vector:
@@
where $\mathbf{f}_{t}\in\mathbb{R}^{m}$ is the SAE feature vector of dimension $m$ for token $t$. The summation method aggregates all token activations, while top-n selects the strongest activations per token.
@@
  • TF-IDF: A classic variation of bag-of-words without neural representations Spärck Jones (1972).
  • Hidden State: Like prior studies Elhage et al. (2024), we did compare to last-token hidden state probing as well.
@@ 4 Classification Tasks, Multimodal Transfer, and Hyperparameter Analysis
We evaluate performance on a suite of open-source classification benchmarks, including binary safety tasks (jailbreak detection, election misinformation, harmful prompt detection) and multi-class scenarios (user intent classification, scenario understanding, abortion intent detection from tweets, banking queries classification). Detailed dataset characteristics are in Appendix A.1.
@@
(b) Token level top-$N$ vs. full binarized features. Token level top-$N$ improves with larger values of $N$, and binarization can worsen this performance. However, binarization of all tokenwise activations reached the best performance of Token level top-$N$ whilst removing the need to compute top-$N$ values, which would be important as $N$ scales, offering a more efficient alternative.
@@ 4.1 Impact of Layer Depth and Model Scale
We extract token-level SAE features and train LR classifiers, comparing the results to TF-IDF and final-layer hidden-state baselines 555we did not benchmark against mean-diff here because that required task to be binary classification.
Figure 3(a) depicts the layer-wise performance for the three model scales across our text-based classification tasks. We observe:
@@
  • Model Scale: Larger models (9B, 9B-IT) achieve consistently higher mean performance (above 0.85 F1) compared to the 2B model. This aligns with the general larger hidden dimension in these models facilitating richer representations.
@@ 4.2 Pooling Strategies and Binarization
We next examine pooling and binarization strategies. Token level max activation pooling methods included no max pooling (top-0), top-20, and top-50 features per token. Binarization is applied after token aggregation.
Figure 3(b) compares two feature selection strategies: (1) no max pooling with summation of all SAE features, and (2) selecting the top-$N$ token level activations (here, 20 and 50), with and without binarization. LR classifiers are trained on the resulting features.
@@
These observations motivate our decision to adopt binarized and no max pooling as a default due to the reduced computational overhead whilst maintaining performance, while acknowledging that token-level top-$N$ might excel for certain tasks.
@@ 5 Multilingual Classification and Transferability
Following the best configurations from previous Section, we extract SAE features from gemma-2-9b and 9b-it (widths of 16K or 131K). We train linear classifiers on one language’s data and test on the same or a different language. We also compare against a simpler SAE feature selection approach, the top-n mean-difference baseline (Mean-Diff), to determine if the entire feature set is necessary.
@@
Figure 4 presents the F1 scores. Pink bars show native SAE training, gold bars show English-trained models tested on other languages, and green bars show English-translated models tested on translated inputs Key findings:
@@
We then examine the robustness of SAE feature extraction with varying amounts of training data.
@@ 6 Behavioral (Action) Prediction
We use the entity-based knowledge conflicts in question answering dataset Longpre et al. (2022), which provides binary correctness labels for model responses. Open-ended generation is performed with vllm Kwon et al. (2023), and answers are scored using inspect ai AI Safety Institute with GPT-4o-mini as the grader.
@@
We focus on predicting 9B-IT’s output correctness. For a given model $M$ (2B, 9B, 9B-IT):
  1. 1.
    We generate open-ended answers to prompts using the model.
  2. 2.
    We use GPT-4o-mini-0718 to label each answer as correct or incorrect.
  3. 3.
    We extract pretrained SAE activations from the input question, with and without provided contexts.
  4. 4.
    We train a logistic regression model to predict the binary correctness label from these extracted features.
We also perform cross-model prediction (e.g., 2B predicting 9B’s correctness), similar to Binder et al. (2024). We fix the SAE width to 16K and compare the quality of predictions using full SAE binarized approach to those using the Top-$N$ mean difference feature method, and analyze auto-interpretable descriptions of features to understand if similar explanations are shared in the top features across models.
Figure 6 summarizes the macro F1 scores across several conditions from the NQ-Swap and inspect_evals datasets: Original context, Question Swap, and Subject Swap. Key findings:
@@
    For multilingual toxicity detection, native training outperforms cross-lingual transfer, though instruction-tuned models (e.g., 9B-IT) may exhibit modestly better transfer as you can see in Appendix A.7 and A.8.
@@ 8 Limitations
Our analysis is restricted to SAEs trained with Jump ReLU activation on Gemma 2 models. This limits generalizability to other model architectures and training paradigms. Future work should explore diverse SAE training strategies and model sources.
@@ Appendix A Appendix
A.1 Models and Dataset Information
@@
A.2 Performance variation on the width
@@
We conduct an analysis of the effect of width scaling on full SAE features among our safety text classification tasks. The evaluation compares models with and without max pooling, as well as binarized and non-binarized activations, to determine their impact on classification performance. Consistently increasing the width results in decline in the mean score across all configurations, with the steepest drop observed in non-binarized cases, which is surprisingly different from Kantamneni et al. (2024) demonstrate the opposite using mean-diff feature SAE selection.
A.3 Multimodal performance
@@
In this study, we also tried the Intel Gemma-2B LLaVA 1.5-based model (Intel/llava-gemma-2b) Hinck et al. (2024) as the foundation for our experiments. For feature extraction, we incorporate pre-trained SAEs from jbloom/Gemma-2b-Residual-Stream-SAEs, trained on the Gemma-1-2B model. These SAEs include 16,384 features (an expansion factor of 8 × 2048) and are designed to capture sparse and interpretable representations.
@@ A.5 Performance Tables
Table 5: Comparison of F1 scores across different layers and top-$N$ token selections. top-$N$ indicates evaluation on the top 10, 20, or 50 mean top-diff SAE features. Original refers to the original input language, while Translated corresponds to translated input to English. Bold values highlight the highest scores for each row.
@@
Gemma2 - 9B Original 0.724 0.764 0.785 0.724 0.764 0.785 0.724 0.764 0.785
Translated 0.723 0.763 0.783 0.723 0.763 0.783 0.723 0.763 0.783
Gemma2 - 9B IT Original 0.722 0.735 0.767 0.722 0.735 0.767 0.722 0.735 0.767
Translated 0.721 0.734 0.764 0.721 0.734 0.764 0.721 0.734 0.764
@@
A.6 Cross Lingual Transfer of Feature Activations
@@
A.6.1 Full SAE learns classifiers find different features than Mean-diff top-N features

### v2

@@ 1 Introduction
Sparse Autoencoders (SAEs) have recently emerged as a promising technique within MI for understanding LLMs. SAEs generally work by learning a compressed, sparse representation of the LLM’s internal activations. This is achieved by up-projecting the dense hidden state of the LLM to a sparser, ideally monosemantic, representation Bricken et al. (2023); Gao et al. (2024b). Identifying semantically meaningful features within LLMs using SAEs allows for deploying these features into explainable classification pipelines. This has the potential to boost performance and detect harmful biases or spurious correlations before they manifest in downstream tasks Bricken et al. (2024). The ability to employ SAE features for classification across diverse settings, ranging from toxicity detection to user intent, offers a scalable form of "model insight" Bowman et al. (2022), which is crucial for building trust, safety, and accountability in high-stakes domains like medicine and law Abdulaal et al. (2024).
Despite the promise of SAEs for MI, surprisingly few systematic studies have provided practical guidance on their use for classification. While promising results have been reported across various tasks Bricken et al. (2024); Kantamneni et al. (2024); Chen et al. (2024), inconsistent experimental protocols, a lack of standardized benchmarks, and limited exploration of key architectural decisions hinder comparability and the development of best practices. Although tools like Transformer Lens Nanda and Bloom (2022) and SAE Lens Joseph Bloom and Chanin (2024) have improved standardization in sampling activations, critical questions about optimal configurations for diverse tasks, particularly in multilingual and multimodal settings, remain unanswered. This makes it challenging to establish the robustness and generalizability of SAE-based classification approaches.
@@ 2.1 Interpretable Feature Extraction
SAEs address these limitations by providing more interpretable feature sets Bricken et al. (2023); Cunningham et al. (2023). Recent scaling efforts have demonstrated SAE viability across LLMs from Claude 3 Sonnet Templeton et al. (2024) to GPT-4 Gao et al. (2024a) with extensions to multimodal architectures like CLIP Bhalla et al. (2024). Although these studies have revealed interpretable feature dimensions and computational circuits Marks et al. (2024); Zhao et al. (2024), they focus mainly on descriptive feature discovery rather than systematic evaluation of their downstream applications. Our work bridges this gap by providing standardized evaluation frameworks for SAE-based classification and cross-modal transfer, establishing quantitative metrics and methods for feature utility across diverse tasks.
@@ 2.2 SAE-Based Classification and its Limitations
Reports have demonstrated that SAE-derived features can outperform traditional hidden-state probing for classification, particularly in scenarios with noisy or limited data with closed datasets Bricken et al. (2024) or simplified tasks Kantamneni et al. (2024). However, more recent studies, such as Wu et al. (2025), suggest that SAEs may not be superior, particularly for model steering (instead of classification). These seemingly conflicting results highlight a critical gap in the current understanding of SAE-based classification: a lack of systematic exploration of how hyperparameters, feature aggregation strategies, and other methodological choices impact performance.
@@ 3 Preliminaries
Experimental Setup Rationale:
Our primary goal is to evaluate pre-trained SAE features for interpretable, zero-shot classification tasks. Accordingly, we selected the Gemma 2 SAE suite as it was the only publicly available family offering matched model backbones (2B, 9B, 9B-IT) with identical training settings and systematic layer and width pairings. We compare against two standard interpretable baselines: linear probes on hidden-state activations and TF-IDF on a bag-of-words representation. We deliberately exclude fine-tuned models, as they operate under a different, less-interpretable paradigm and fall outside our zero-shot evaluation scope. The TF-IDF baseline serves as a strong, classic non-neural benchmark for interpretability and performance.
@@
Since SAEs generate token-level feature activations, an essential step in classification is aggregating these activations into a fixed-size sequence representation. Without pooling, the model lacks a structured way to combine token-level representations. Previous NLP works have explored various pooling strategies for feature aggregation in neural representations Shen et al. (2018). However, it remains unclear which pooling method is most effective for LLMs’ SAE features. We systematically evaluate different pooling approaches (displayed in 2, considering (1) Top-$N$ feature selection per token  [^3] Token-level top-$N$ where n=0 indicates the absence of max pooling. Karvonen et al. (2025) and (2) summation-based aggregation444this approach is also adopted by parallel research Brinkmann et al. (2025). which collapses token-level activations into a single sequence vector:
@@
where $\mathbf{f}_{t}\in\mathbb{R}^{m}$ is the SAE feature vector of dimension $m$ for token $t$. The summation method aggregates all token activations, while top-n selects the strongest activations per token. Further details are provided in A.1.
@@
  • TF-IDF: Classic bag-of-words variation without neural representations Spärck Jones (1972).
  • Hidden State: Like prior studies features_as_classifiers, we did compare to last-token hidden state probing as well.
@@ 4 Classification Tasks, Multimodal Transfer, and Hyperparameter Analysis
We targeted scalable, safety-relevant binary classification tasks—jailbreak detection, harmful-prompt screening, and multilingual toxicity—to stress-test generality while keeping evaluation simple and comparable. Concretely, we selected publicly available datasets drawn from MTEB and other widely used classification corpora to ensure reproducibility and sufficient scale Muennighoff et al. (2023). We prioritized (i) clear binary labels, (ii) coverage across multiple languages, and (iii) permissive licensing. At the time of experimentation, the pool of multilingual binary datasets was limited, so we focused on these three tasks; broadening the task set is an important direction for future work. Detailed dataset characteristics are in Appendix A.2.
@@
(b) Token level top-$N$ vs. full binarized features. Token level top-$N$ improves with larger values of NN, and binarization can worsen this performance. However, binarization of all tokenwise activations reached the best performance of Token level top-$N$ whilst removing the need to compute top-NN values, which would be important as NN scales, offering a more efficient alternative.
@@ 4.1 Impact of Layer Depth and Model Scale
We extract token-level SAE features and train LR classifiers, comparing the results to TF-IDF and final-layer hidden-state baselines 555we did not benchmark against mean-diff here because that required task to be binary classification. Figure 3(a) depicts the layer-wise performance for the three model scales across our text-based classification tasks. We observe:
@@
  • Model Scale: Larger models (9B, 9B-IT) achieve consistently higher mean performance (above 0.85 F1) compared to the 2B model. This aligns with larger hidden dimension in these models having richer representations.
@@ 4.2 Pooling Strategies and Binarization
We next examine pooling and binarization strategies. Token level max activation pooling methods included no max pooling (top-0), top-20, and top-50 features per token. Binarization is applied after token aggregation. Figure 3(b) compares two feature selection strategies: (1) no max pooling with summation of all SAE features, and (2) selecting the top-$N$ token level activations (here, 20 and 50), with and without binarization. LR classifiers are trained on the resulting features with L2 regularization.
@@
These observations motivate our decision to adopt binarized and no max pooling as a default due to theoretical reduced computational overhead whilst maintaining performance, while acknowledging that token-level top-$N$ might excel for certain tasks.
@@ 5 Multilingual Classification and Transferability
Following the best configurations from previous Section, we extract SAE features from gemma-2-9b and 9b-it (widths of 16K or 131K). We train linear classifiers on one language’s data and test on the same or a different language. We also compare against a simpler SAE feature selection approach, the top-n mean-difference baseline (Mean-Diff) Kantamneni et al. (2024), to determine if the entire feature set is necessary.
@@
Figure 4 presents the F1 scores. Pink bars show native SAE training, gold bars show English-trained models tested on other languages, and green bars show English-translated models tested on translated inputs:
@@
However, when considering average rather than peak performance, Mean-Diff top-N selection actually outperforms SAE features, providing a higher mean F1 score and lower variance. This suggests it may be preferable in scenarios where stability across tasks is prioritized over peak performance. We then examine the robustness of SAE feature extraction with varying amounts of training data.
@@
Clarifying differences from Kantamneni et al. (2024)
The use of L1 sparsity methods to perform feature selection, mean-difference approach of Kantamneni et al. (2024), demonstrates strong performance and that a small number of features can contain most of the task-relevant information. However, for the specific task of our multilingual toxicity detection, the aggregated binarisation method from all features appears to preserve a stronger signal and greater transferability across languages in native and translated settings. This is in contrast to Kantamneni et al. (2025), and therefore, future work is needed to clarify the task sensitivity of the divergent findings. Major differences on feature selection methods may also drive differences and future work will focus on understanding the impact of different methods on varied interpretability approaches.
@@ 6 Behavioral (Action) Prediction
We use the entity-based knowledge conflicts in question answering dataset Longpre et al. (2022), which provides binary correctness labels for model responses. Open-ended generation is performed with vllm Kwon et al. (2023), and answers are scored using inspect ai AI Safety Institute (2024) with GPT-4o-mini as the grader.
@@
We focus on predicting 9B-IT’s output correctness. For a given model $M$ (2B, 9B, 9B-IT): 1) We generate open-ended answers to prompts using the model. 2) We use GPT-4o-mini-0718 to label each answer as correct or incorrect. 3) We extract pretrained SAE activations from the input question, with and without provided contexts. 4) We train a logistic regression model to predict the binary correctness label from these extracted features.
We also perform cross-model prediction (e.g., 2B predicting 9B’s correctness), similar to Binder et al. (2024). We fix the SAE width to 16K and compare the quality of predictions using full SAE binarized approach to those using the Top-$N$ mean difference feature method, and analyze auto-interpretable descriptions of features to understand if similar explanations are shared in the top features across models. Figure 6 summarizes the macro F1 scores across several conditions from the NQ-Swap and inspect_evals datasets: Original context, Question Swap, and Subject Swap. Key findings:
@@
    For multilingual toxicity detection, native training outperforms cross-lingual transfer, though instruction-tuned models (e.g., 9B-IT) may exhibit modestly better transfer as you can see in Appendix A.8 and A.9.
@@ 8 Limitations
Our primary analysis is restricted to SAEs trained with Jump ReLU activation on Gemma 2 models as they were the only open-source models available that provided SAE’s across varying layers, widths, and model sizes. This could potentially limit generalizability to other model architectures and training paradigms. Future work should explore diverse SAE training strategies and model sources.
@@
Future Work: Robustness under Domain Shift
A crucial next step is to investigate how SAE-derived features behave when the input distribution changes. This includes examining covariate, subpopulation, and temporal shifts by training probes on one domain and evaluating on held-out domains (e.g., news→social media; formal→informal), measuring activation drift and the stability of feature–label associations. This evaluation will clarify whether the observed transferability reflects domain-agnostic structure or domain-specific correlations.
Beyond robustness, there is a need to expand the task set beyond safety-oriented binary classification to include multilabel and non-safety tasks and additional multilingual benchmarks.
@@ Appendix A Appendix
A.1 Details on Pooling Methods
Top-NN Feature Selection per Token
In our approach, the top-$N$ feature selection per token is performed as follows:
Step 1
For each token $t$ in a sequence, we consider its corresponding SAE activation vector:
$f_{t}\in\mathbb{R}^{m},\quad t=1,2,\dots,n,$
where $m$ is the SAE dimension and $n$ is the sequence length.
Step 2
For each token-level activation vector $f_{t}$, we keep only the top $N$ largest activation values, setting all other activations to zero:
[…]
Step 3
We then aggregate these sparse vectors across all tokens by summation to obtain a fixed-size sequence-level representation $F$:
$F=\sum_{t=1}^{n}\tilde{f}_{t}.$
Thus, the selection is performed per token independently (not across the entire dataset at once). This ensures each token contributes its most salient features, and then we aggregate token-level sparse activations into a sequence-level vector.
Top-NN Mean-Difference Selection
The top-$N$ mean-difference selection method is a supervised feature selection approach performed at the dataset level, as follows:
Step 1
For each SAE dimension $i$, compute the absolute difference between the mean activation for the positive class $C^{+}$ and the negative class $C^{-}$ over the entire training set:
$d_{i}\;=\;\left|\frac{1}{\lvert C^{+}\rvert}\sum_{x\in C^{+}}F_{x}(i)\;-\;\frac{1}{\lvert C^{-}\rvert}\sum_{x\in C^{-}}F_{x}(i)\right|,$
where $F_{x}(i)$ is the aggregated activation of dimension $i$ for instance $x$.
Step 2
Select the top $N$ SAE dimensions with the largest $d_{i}$ values. This selection is done once at the dataset level using the training data.
Step 3
For subsequent classification, keep only these top $N$ dimensions for all instances.
In other words, the mean-difference selection is computed using activations aggregated across all tokens and all instances in the training dataset to identify globally discriminative SAE dimensions.
A.2 Models and Dataset Information
@@
A.3 Performance variation on the width
@@
We conduct an analysis of the effect of width scaling on full SAE features among our safety text classification tasks. The evaluation compares models with and without max pooling, as well as binarized and non-binarized activations, to determine their impact on classification performance. Consistently increasing the width results in decline in the mean score across all configurations, with the steepest drop observed in non-binarized cases, which is surprisingly different from Kantamneni et al. (2024) demonstrate the opposite using mean-diff feature SAE selection. A complete table of our results across variations are available at the following anonymous link https://docs.google.com/spreadsheets/d/1zUTXBdsorzthBLwMUoXNBP-X5lrUysnNL0iLYdBZ1HU/edit?usp=sharing.
A.4 Multimodal performance
@@
In this study, we also tried the Intel Gemma-2B LLaVA 1.5-based model (Intel/llava-gemma-2b) Hinck et al. (2024) as the foundation for our experiments. For feature extraction, we incorporate pre-trained SAEs from jbloom/Gemma-2b-Residual-Stream-SAEs (RELU-based), trained on the Gemma-1-2B model. These SAEs include 16,384 features (an expansion factor of 8 × 2048) and are designed to capture sparse and interpretable representations.
@@ A.6 Performance Tables
Table 5: Comparison of F1 scores across different layers and top-$N$ token selections. top-NN indicates evaluation on the top 10, 20, or 50 mean top-diff SAE features. Original refers to the original input language, while Translated corresponds to translated input to English. Bold values highlight the highest scores for each row.
@@
Gemma2 - 9B Original 0.72 0.76 0.79 0.72 0.76 0.79 0.72 0.76 0.79
Translated 0.72 0.76 0.78 0.72 0.76 0.78 0.72 0.76 0.78
Gemma2 - 9B IT Original 0.72 0.74 0.77 0.72 0.74 0.77 0.72 0.74 0.77
Translated 0.72 0.73 0.76 0.72 0.73 0.76 0.72 0.73 0.76
@@
A.7 Cross Lingual Transfer of Feature Activations
@@
A.7.1 Full SAE learns classifiers find different features than Mean-diff top-N features